# Covering in Hamming and Grassmann Spaces: New Bounds and Reed–Solomon-Based Constructions

### v1

@@ IV-E Coverage Fraction
where $A_{w}$ is the number of codewords $c\in{\mathcal{C}}$ of weight $w$, and $I(w,\tau):=|B(c_{1},\tau)\cap B(c_{2},\tau)|$ for $c_{1},c_{2}\in{\mathcal{C}}$ with $d_{{\rm H}}(c_{1},c_{2})=w$. Now, (43) is known to hold for $d\leq w\leq n$ (e.g., see [Macwilliams77]), and (44) follows by a counting argument, where $u$ (resp. $v$) represents the number of indices in $\{1,\dots,n\}$ where $y$ and $c_{1}$ (resp. $c_{2}$) agree, and $z$ represents the number of indices where $c_{1}$, $c_{2}$ and $y$ agree. Substituting $|{\mathcal{C}}|=q^{k}=q^{n-d+1}$ into (45) gives (42). ∎
@@ V Character-Reed–Solomon Code
Recall from Definition 3 that the message space ${\mathcal{F}}_{p}(k,q)$ of CP codes consists of polynomials $f\in{\mathcal{F}}(k,q)$ with $f_{jp}=0$ for all integers $j\geq 0$. This condition was used to establish minimum distance properties of CP codes [Hessam22]. Since our focus in this work is on covering rather than packing, we utilize the full message space ${\mathcal{F}}(k,q)$ and generalize the definition of CP codes below to what we call character-Reed–Solomon (CRS) codes.
@@ V-B Covering Algorithm
As mentioned in the introduction, CP codes have recently been used for the quantization problem in MISO systems with limited feedback [gooty2025precodingdesignlimitedfeedbackmiso], and the quantization error was characterized using the covering radius of CP codes. Moreover, a decoding algorithm for CP codes over prime fields was recently proposed [riasat2024decodinganalogsubspacecodes]. However, as discussed at the beginning of Section IV, decoding algorithms do not produce good coverage due to the decoding radius being much smaller than the covering radius. Therefore, we propose Algorithm 2 below for covering lines in $G_{1,n}({\mathbb{C}})$ using CRS codes. To describe the idea behind the algorithm, consider the map $\psi_{\beta}$ defined in (58), which can be regarded as an extension of the 11-to-$q/p$ map $\chi_{\beta}^{-1}$ to all of $\mathbb{C}$}, obtained by first mapping any $z\in{\mathbb{C}}$ to its closest point (in euclidean distance) in $\chi_{\beta}({\mathbb{F}}_{q})$ then taking its pre-image:
@@
In a practical implementation of Algorithm 2, one may avoid Step 8 altogether by modifying $\operatorname{GRS-cover}$ in Algorithm 1 to directly return the message polynomial in Step 7 rather than its re-encoding. Moreover, the pre-images $u_{j}\in\psi_{\beta}(y_{j})$ in Step 5 under the character map $\chi_{\beta}$ may be chosen using a “best-of-NN” approach for some fixed $N$. In other words, out of $N$ random assignments of the elements of $\mathbb{F}_{q}$ to the $u_{i}$ under $\psi_{\beta}$, the one giving the smallest chordal distance of the $\operatorname{CRS}$ codeword from $y$ is chosen. This is because, in general, each point in $\chi_{\beta}(\mathbb{F}_{q})$ has $q/p$ distinct pre-images in $\mathbb{F}_{q}$, so a brute-force approach would need to test $(q/p)^{n}$ assignments to find the optimal one.
@@ V-C Improved Upper Bound on Average Covering Radius
In other words, $\check{y}_{j}$ is the closest point to $y_{j}$ in $\Psi$ (in Euclidean distance). The following theorem improves the upper bound on the average chordal covering radius of $\operatorname{CP}_{n,\beta}({\mathcal{F}}_{p}(k,p))=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,p))$ by Gooty et al. [gooty2025precodingdesignlimitedfeedbackmiso, Theorem 1].
@@
The first expression inside the $\min$ in (61) is the square root of the bound on the mean-squared quantization error given by Gooty et al. [gooty2025precodingdesignlimitedfeedbackmiso, Theorem 1]. To obtain the improved bound, let $f\in{\mathcal{F}}(k,p)$ denote the output of $\operatorname{GRS-decode}$ in Step 8 of $\operatorname{CRS-cover}({\mathcal{C}},y)$. Then, by the triangle inequality,
@@
Theorem 8 provides sufficient SNR and rate conditions under which CRS codes achieve a small average covering radius. Comparing this bound with the random-coding expectation of Theorem 2 yields the following corollary, which shows that CRS codes asymptotically achieve the average covering radius of random codes at high rate and high SNR.
@@
Consider a distribution $\mathcal{D}$} on ${\mathbb{C}}^{n}$, where the amplitudes are i.i.d. with mean $\mu$ and variance $\sigma^{2}{}$, and are independent of the phases. Then, over $\mathcal{D}$}, the average covering radius of a $\operatorname{CRS}$ code of rate $R$ over ${\mathbb{F}}_{p}$ such that
$\frac{\mu^{2}}{\mu^{2}+\sigma^{2}}\geq\max\Bigg\{$ $\frac{1-c}{1-c+2n(cR+R-1)},$ (70)
$\frac{1}{(cR+R-1)^{2}}$ $\Bigg(1-\frac{p^{-R(1+\frac{1}{n-1})}\Gamma(1+\tfrac{1}{2(n-1)})^{2}}{(1+\frac{1}{2(n-1)p^{nR}})^{2}}\Bigg)\Bigg\},$
where $c:=\sqrt{\cos(2\pi/p)}$ and $\Gamma$ is the gamma function, is at most the average covering radius of a random code of size $p^{nR}$ in $G_{1,n}({\mathbb{C}})$. In other words, for sufficiently large $n$, $\operatorname{CRS}$ codes asymptotically achieve the average covering radius of random codes under the following SNR bound provided $R>(1+\sqrt{1-p^{-R}})/(c+1)$:
$\dfrac{\mu^{2}}{\sigma^{2}}$ $\geq\frac{1-p^{-R}}{(cR+R-1)^{2}-(1-p^{-R})}.$ (71)
In particular, for $R\to 1$, the above conclusion holds under:
$\dfrac{\mu^{2}}{\sigma^{2}}$ $\geq p-1+2\pi^{2}.$ (72)
@@
By Wendel’s inequality [wendel48],
$\frac{\Gamma(M+1)}{\Gamma(a_{n}+M-1)}$ $\geq\frac{M^{3-a_{n}}}{a_{n}+M-2}=\frac{p^{-\frac{R}{2}(1+\frac{1}{n-1})}}{1+\frac{p^{-nR}}{2(n-1)}},$ (75)
since $a_{n}-2\in(0,1)$ by (74). Therefore, by Theorem 8 and Eqs. 73, 74 and 75, CRS codes achieve the average covering radius of random codes whenever (70) holds.
Finally, note that as $n\to\infty$,
$\frac{p^{-R(1+\frac{1}{n-1})}\Gamma(1+\tfrac{1}{2(n-1)})^{2}}{(1+\frac{1}{2(n-1)p^{nR}})^{2}}$ $\sim p^{-R},$ (76)
where the asymptotic notation $A(n)\sim B(n)$ conventionally means that $A(n)/B(n)\to 1$. Thus, (71) follows by substituting (76) into (70) provided that $R>(1+\sqrt{1-p^{-R}})/(c+1)$, and (72) follows by observing that the right-hand side of (71) is asymptotic to $p-1+2\pi^{2}$ for $R\to 1$. ∎
@@ V-D Simulation Results and Comparison with Random Coding Bound
Figure 6 shows the Monte Carlo simulated average covering radius of ${\mathcal{C}}=\operatorname{CRS}_{6,1}({\mathcal{F}}(k,7))$ for $1\leq k<6$ computed using Algorithm 2 with $\operatorname{GRS-decode}\in\{\hbox{BW},\hbox{GS}\}$, together with theoretical bounds (compare this with Figure 1). The input vector $y\in{\mathbb{C}}^{n}$ for each simulation was chosen to have i.i.d. coordinates drawn from a complex normal distribution so that the amplitudes follow a Rayleigh distribution with mean $1$. The average chordal distance from $y$ to the output codeword was computed over $500$ data points for each $k$. The random coding bound is given by (23) in Theorem 2. The Gooty et al. upper bound is given by the square root of their bound on the mean squared quantization error [gooty2025precodingdesignlimitedfeedbackmiso, Theorem 1], which is improved upon by Theorem 8.
11223344550.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911Dimension ($k$)Average Covering RadiusUpper Bound [gooty2025precodingdesignlimitedfeedbackmiso] Upper Bound (Theorem 8) Algorithm 2 (BW)Algorithm 2 (GS)Random Codes (Theorem 2)
@@
Figure 7 compares the upper bound on the average covering radius of an $(n,k)_{p}$ CRS code and the random coding bound in the extreme high-rate regime $k/n\to 1$, where n=p1n=p-1 and k=n1k=n-1, for primes $p\leq 70571$. The bounds are evaluated under a source distribution with mean $\mu=n$ and variance $\sigma^{2}=1$. As $p\to\infty$, both the random-coding bound and the upper bound derived in Theorem 8 vanish, while the ratio between them remains bounded by a constant. In particular, the average covering radius of the CRS code remains sandwiched between the two bounds and within a fixed constant factor of the random-coding bound, in accordance with Corollary 3.
15,00015{,}00030,00030{,}00045,00045{,}00060,00060{,}00000.020.020.040.040.060.060.080.080.10.1Field Size ($p$)Average Covering RadiusUpper Bound (Theorem 8)Random Codes (Theorem 2)
Figure 7: Convergence of upper and lower bounds on the average covering radius of a $(p-1,p-2)_{p}$ CRS code for prime $p$
@@ VI Conclusion and Future Directions
From an information-theoretic perspective, we introduced the notion of the average covering radius for both block and subspace codes as a meaningful measure of average distortion in quantization problems. Using tools from one-shot compression and rate–distortion theory, we derived explicit random-coding bounds on the average covering radius in both Hamming and Grassmann spaces. These bounds provide fundamental benchmarks against which structured code constructions can be evaluated. In particular, our numerical results demonstrated that the RS-based constructions outperform random codebooks in various fixed block-length scenarios. The analysis provides new insights into the role of algebraic structures in covering problems and applications.

### v2

@@ IV-E Coverage Fraction
where $A_{w}$ is the number of codewords $c\in{\mathcal{C}}$ of weight $w$, and $I(w,\tau):=|B(c_{1},\tau)\cap B(c_{2},\tau)|$ for $c_{1},c_{2}\in{\mathcal{C}}$ with $d_{{\rm H}}(c_{1},c_{2})=w$. Now, (43) is known to hold for $d\leq w\leq n$ (e.g., see [20]), and (44) follows by a counting argument, where $u$ (resp. $v$) represents the number of indices in $\{1,\dots,n\}$ where $y$ and $c_{1}$ (resp. $c_{2}$) agree, and $z$ represents the number of indices where $c_{1}$, $c_{2}$ and $y$ agree. Substituting $|{\mathcal{C}}|=q^{k}=q^{n-d+1}$ into (45) gives (42). ∎
@@ V Character-Reed–Solomon Code
Recall from Definition 3 that the message space ${\mathcal{F}}_{p}(k,q)$ of CP codes consists of polynomials $f\in{\mathcal{F}}(k,q)$ with $f_{jp}=0$ for all integers $j\geq 0$. This condition was used to establish minimum distance properties of CP codes [36]. Since our focus in this work is on covering rather than packing, we utilize the full message space ${\mathcal{F}}(k,q)$ and generalize the definition of CP codes below to what we call character-Reed–Solomon (CRS) codes.
@@ V-B Covering Algorithm
As mentioned in the introduction, CP codes have recently been used for the quantization problem in MISO systems with limited feedback [14], and the quantization error was characterized using the covering radius of CP codes. Moreover, a decoding algorithm for CP codes over prime fields was recently proposed [28]. However, as discussed at the beginning of Section IV, decoding algorithms do not produce good coverage due to the decoding radius being much smaller than the covering radius. Therefore, we propose Algorithm 2 below for covering lines in $G_{1,n}({\mathbb{C}})$ using CRS codes. To describe the idea behind the algorithm, consider the map $\psi_{\beta}$ defined in (58), which can be regarded as an extension of the 11-to-$q/p$ map $\chi_{\beta}^{-1}$ to all of $\mathbb{C}$}, obtained by first mapping any $z\in{\mathbb{C}}$ to its closest point (in euclidean distance) in $\chi_{\beta}({\mathbb{F}}_{q})$ then taking its pre-image:
@@
In a practical implementation of Algorithm 2, one may avoid Step 8 altogether by modifying $\operatorname{GRS-cover}$ in Algorithm 1 to directly return the message polynomial rather than its re-encoding in Step 7. Moreover, the pre-images $u_{j}\in\psi_{\beta}(y_{j})$ in Step 5 under the character map $\chi_{\beta}$ may be chosen using a “best-of-NN” approach for some fixed $N$. In other words, out of $N$ random assignments of the elements of $\mathbb{F}_{q}$ to the $u_{i}$ under $\psi_{\beta}$, the one giving the smallest chordal distance of the $\operatorname{CRS}$ codeword from $y$ is chosen. This is because, in general, each point in $\chi_{\beta}(\mathbb{F}_{q})$ has $q/p$ distinct pre-images in $\mathbb{F}_{q}$, so a brute-force approach would need to test $(q/p)^{n}$ assignments to find the optimal one.
@@ V-C Improved Upper Bound on Average Covering Radius
In other words, $\check{y}_{j}$ is the closest point to $y_{j}$ in $\Psi$ (in Euclidean distance). The following theorem improves the upper bound on the average chordal covering radius of $\operatorname{CP}_{n,\beta}({\mathcal{F}}_{p}(k,p))=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,p))$ by Gooty et al. [14, Theorem 1].
@@
The first expression inside the $\min$ in (61) is the square root of the bound on the mean-squared quantization error given by Gooty et al. [14, Theorem 1]. To obtain the improved bound, let $f\in{\mathcal{F}}(k,p)$ denote the output of $\operatorname{GRS-decode}$ in Step 8 of $\operatorname{CRS-cover}({\mathcal{C}},y)$. Then, by the triangle inequality,
@@
Theorem 8 provides sufficient SNR and rate conditions under which CRS codes achieve a small average covering radius. Comparing this bound with the random-coding expectation of Theorem 2 yields the following corollary, which shows that CRS codes achieve average covering radii within a constant factor of random codes at high rate and high SNR.
@@
Consider a distribution $\mathcal{D}$} on ${\mathbb{C}}^{n}$, where the amplitudes are i.i.d. with mean $\mu$ and variance $\sigma^{2}{}$, and independent of the phases, such that $\mu^{2}/\sigma^{2}\to\infty$ as $n\to\infty$. Then, for sufficiently large $n$, the average covering radius over $\mathcal{D}$} of a $\operatorname{CRS}$ code of rate $R>1-1/(\ln p)$ over ${\mathbb{F}}_{p}$ is within a factor of $2\sqrt{p(1-R)}$ of the average covering radius of a random code of size $p^{nR}$ in $G_{1,n}({\mathbb{C}})$.
@@
Using Stirling’s approximation, as $n\to\infty$,
$\frac{\Gamma(M+1)}{\Gamma(a_{n}+M-1)}$ $\sim\frac{M^{M+1/2}e^{-M}}{(a_{n}+M-2)^{a_{n}+M-3/2}e^{-(a_{n}+M-2)}}$
$=e^{a_{n}-2}\frac{M^{M+1/2}}{(a_{n}+M-2)^{a_{n}+M-3/2}}$
$=e^{a_{n}-2}\frac{M^{M+1/2-a_{n}-M+3/2}}{(\frac{a_{n}+M-2}{M})^{a_{n}+M-3/2}}$
$=\left(\frac{e}{M}\right)^{a_{n}-2}\frac{1}{(1+\frac{a_{n}-2}{M})^{a_{n}+M-3/2}}$
$\sim\left(\frac{e}{M}\right)^{a_{n}-2}\frac{1}{e^{a_{n}-2}(1+\frac{a_{n}-2}{M})^{a_{n}-5/2}}$
$\sim M^{2-a_{n}}$
$=M^{-\frac{1}{2(n-1)}}$
$\sim p^{-\frac{R}{2}},$ (72)
and, by Theorem 8,
$\overline{\rho}_{{\rm ch}}({\mathcal{C}})$ $\leq\sqrt{1-\frac{(cR+R-1)^{2}\mu^{2}}{\mu^{2}+\sigma^{2}}}$
$\sim\sqrt{1-(2R-1)^{2}},$ (73)
where ${\mathcal{C}}=\operatorname{CRS}_{n}({\mathcal{F}}_{p}(k,p))$. Here, the asymptotic notation $A(n)\sim B(n)$ conventionally means that $A(n)/B(n)\to 1$ as $n\to\infty$. Letting $R=1-\varepsilon$ for $\varepsilon<1/(\ln p)$, we have
$p^{-\frac{R}{2}}$ $=\sqrt{p^{-1}e^{\varepsilon\ln p}}\sim\sqrt{\frac{1+\varepsilon\ln p}{p}},$ (74)
and
$\sqrt{1-(2R-1)^{2}}$ $=\sqrt{1-(1-2\varepsilon)^{2}}\sim 2\sqrt{\varepsilon},$ (75)
i.e., the ratio of the average covering radii of a random code and $\mathcal{C}$}, by Eqs. 70, 71, 72, 73, 74 and 75, is asymptotically at least
$\frac{p^{-\frac{R}{2}}}{\sqrt{1-(2R-1)^{2}}}$ $\sim\frac{1}{2\sqrt{\varepsilon p}}.$ (76)
Therefore, $\mathcal{C}$} achieves the average covering radius of random codes within a factor of $2\sqrt{\varepsilon p}$ for sufficiently large $n$. ∎
Note that $2\sqrt{p(1-R)}\sim 2\sqrt{n-k}{}$ whenever $p=n+o(n)$. Classical results from analytic number theory (see, e.g., [1]) guarantee the existence of such $p$ for any $n$.
@@ V-D Simulation Results and Comparison with Random Coding Bound
Figure 6 shows the Monte Carlo simulated average covering radius of ${\mathcal{C}}=\operatorname{CRS}_{6,1}({\mathcal{F}}(k,7))$ for $1\leq k<6$ computed using Algorithm 2 with $\operatorname{GRS-decode}\in\{\hbox{BW},\hbox{GS}\}$, together with theoretical bounds (compare this with Figure 1). The input vector $y\in{\mathbb{C}}^{n}$ for each simulation was chosen to have i.i.d. coordinates drawn from a complex normal distribution so that the amplitudes follow a Rayleigh distribution with mean $1$. The average chordal distance from $y$ to the output codeword was computed over $500$ data points for each $k$. The random coding bound is given by (23) in Theorem 2. The Gooty et al. upper bound is given by the square root of their bound on the mean squared quantization error [14, Theorem 1], which is improved upon by Theorem 8.
11223344550.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911Dimension ($k$)Average Covering RadiusUpper Bound [14] Upper Bound (Theorem 8) Algorithm 2 (BW)Algorithm 2 (GS)Random Codes (Theorem 2)
@@
Figure 7 compares the upper bound on the average covering radius of an $(n,k)_{p}$ CRS code and the random coding bound in the extreme high-rate regime $k/n\to 1$, where k=n1k=n-1 and $p$ is the smallest prime >n>n, for $n\leq 10^{4}$. The bounds are evaluated under a source distribution with mean $\mu=n$ and variance $\sigma^{2}=1$. As $n\to\infty$, both the random-coding bound and the upper bound derived in Theorem 8 vanish, while the ratio between them approaches $2$. In particular, the average covering radius of the CRS code remains sandwiched between the two bounds and within a fixed constant factor of the random-coding bound, in accordance with Corollary 3.
1,0001{,}0002,0002{,}0003,0003{,}0004,0004{,}0005,0005{,}0006,0006{,}0007,0007{,}0008,0008{,}0009,0009{,}00000.020.020.040.040.060.060.080.08Length ($n$)Average Covering RadiusUpper Bound (Theorem 8)Random Codes (Theorem 2)
Figure 7: Bounds on the average covering radius of an $(n,n-1)_{p}$ CRS code for $p=n+o(n)$
@@ VI Conclusion and Future Directions
From an information-theoretic perspective, we introduced the notion of the average covering radius for both block and subspace codes as a meaningful measure of average distortion in quantization problems. Using tools from one-shot compression and rate-distortion theory, we derived explicit random-coding bounds on the average covering radius in both Hamming and Grassmann spaces. These bounds provide fundamental benchmarks against which structured code constructions can be evaluated. In particular, our numerical results demonstrated that the RS-based constructions outperform random codebooks in various fixed block-length scenarios. The analysis provides new insights into the role of algebraic structures in covering problems and applications.